# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 7.1. Semicoideals of block subspaces
We have that $A_{G}\leq^{*}B$, so let $a\in E^{[<\infty]}\mathord{\upharpoonright}A_{G}\cap E^{[<\infty]}\cap B$ be such that $A_{G}/a\leq B/a\leq A_{a}$. Therefore, $(a,B)$ captures $\mathcal{D}$ in the ground model, so $(a,B)$ captures $\mathcal{D}$ in the generic extension (see the remark after Definition 7.7). Since $A_{G}\in[a,B]$, we have that $(r_{n}(A_{G}),B)\in\mathcal{D}$ for some $n$, as desired. ∎
@@
Recall that for an uncountable cardinal $\kappa$, the forcing poset $\operatorname{Coll}(\omega,<\!\kappa)$ is the Lévy collapse, which collapses $\kappa$ to $\omega_{1}$. We shall follow the proof layout presented in [NN18] (for the proof of Theorem 16). The following lemma is due to Solovay, and a proof may be found in [solovay]. For the rest of this section, we fix a ground model $\mathbf{V}$ of $\mathsf{ZFC}$, and all vector spaces are over $\mathbb{F}_{2}$.
@@
Let $\mathcal{H}\in\mathbf{V}[G]$ be a selective coideal of $E^{[\infty]}$, and let $\mathcal{X}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ be a subset of $E^{[\infty]}$. Since every set in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is ordinal-definable from some real in $\mathbf{V}[G]$, there exists a ternary formula $\phi$, some real $r\in\mathbf{V}[G]$ and a sequence of ordinals $\alpha$ such that for all $B\in(E^{[\infty]})^{\mathbf{V}[G]}$:
@@
We need to show that for all $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,B]\subseteq\mathcal{X}$ or $a,B]\subseteq\mathcal{X}$. We fix some $A\in\mathcal{H}$.
Since $\kappa$ is Mahlo, we may fix some inaccessible $\lambda<\kappa$ such that $\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]\in[G\mathord{\upharpoonright}\lambda]$, and $\mathbf{V}[G\mathord{\upharpoonright}\lambda]\models\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$ is a selective coideal, and denote $\overline{\mathcal{H}}:=\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. By the $\kappa$-chain condition of $\operatorname{Coll}(\omega,<\!\kappa)$, we may also assume that $A\in\overline{\mathcal{H}}$ and $r\in\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. Let $\phi^{*}$ be the formula such that $\phi^{*}(x,y,w)$ holds iff $\Vdash_{\operatorname{Coll}(\omega,<\kappa)}\phi(\dot{x},\check{y},\check{w})$. Let:
$\overline{\mathcal{X}}:=\{B\in E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]:V[G\mathord{\upharpoonright}\lambda]\models\phi[B,r,\alpha]\}.$
@@
$B\in\overline{\mathcal{X}}$ $\iff\mathbf{V}[G]\models\phi[B,r,\alpha]$
@@
Let $\mathcal{U}$ be a $(\overline{\mathcal{H}},\leq^{*})$-generic filter such that $A\in\mathcal{U}$. Note that by Proposition 7.6, $\mathcal{U}$ is a selective ultrafilter. Since $(\overline{\mathcal{H}},\leq^{*})$ adds no reals, we have that $E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]=E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. This implies that $\overline{\mathcal{X}}=\mathcal{X}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Furthermore, we have that:
@@
Let $\mathbb{P}\in\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$ be some forcing notion which adds an element of $E^{[\infty]}$, and let $B\in E^{[\infty]}\cap\mathbf{V}[G][\mathcal{U}]$ be $\mathbb{P}$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Then:
@@
We work in $\mathbf{V}[G][\mathcal{U}]$. Since $\kappa$ is inaccessible, $\operatorname{Coll}(\omega,<\!\lambda)*\dot{\mathbb{P}}$ is a poset in $\mathbf{V}$ of size $\kappa$, so by Lemma 7.10 applied to the model $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$ there exists a filter $G^{*}$, $\operatorname{Coll}(\omega,<\!\kappa)$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]$ such that:
@@
We may now prove Corollary 1.8. Given an almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$, we let:
$\overline{\mathcal{A}}:=\{B\in E^{[\infty]}:B\leq A\text{ for some }A\in\mathcal{A}\}.$
@@
Suppose for a contradiction that $\mathcal{A}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is a full mad family of block subspaces. Note that $\overline{\mathcal{A}}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$. By Proposition 4.13, $\mathcal{H}(\mathcal{A})$ is a selective coideal, so by Theorem 1.7 we have that $\mathcal{A}$ is $\mathcal{H}$-Ramsey. Let $B\in\mathcal{H}(\mathcal{A})$, and let $C\in\mathcal{H}(\mathcal{A})\mathord{\upharpoonright}A$ be such that $\emptyset,C]\subseteq\overline{\mathcal{A}}$ or $\emptyset,C]\subseteq\overline{\mathcal{A}}^{c}$. Since $C\in\mathcal{H}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})\cap\overline{\mathcal{A}}=\emptyset$, the first case is not possible. On the other hand, since $\mathcal{A}$ is mad, there exists some $A\in\mathcal{A}$ which is compatible with $C$, i.e. there exists some $D\in E^{[\infty]}$ such that $D\leq C$ and $D\leq A$. Then $D\in[\emptyset,C]\cap\overline{\mathcal{A}}$, another contradiction. ∎
@@ 8. Further remarks and open problems
Most of the current approaches to the non-existence of mad families of $\omega]^{\omega}$ employ some Ramsey theoretic techniques (such as [M77] [NN18] [ST19]). Consequently, it is difficult to use these approaches to prove the non-existence of mad families of block subspaces over $\mathbb{F}$ for $|\mathbb{F}|>2$. Perhaps the approach highlighted in [T18] may be used to answer the following question:
@@
Let $\kappa$ be an inaccessible cardinal, and let $G$ be $\operatorname{Coll}(\omega,<\!\kappa)$-generic. Are there mad families of block subspaces (over any countable field) in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$? What if $\kappa$ is Mahlo?

### v5

@@ 7.1. Semicoideals of block subspaces
We have that $D\leq^{*}B$, so let $a\in E^{[<\infty]}\mathord{\upharpoonright}D\cap E^{[<\infty]}\cap B$ be such that $D/a\leq B/a\leq A_{a}$. Therefore, $(a,B)$ captures $\mathcal{D}$ in the ground model, so $(a,B)$ captures $\mathcal{D}$ in the generic extension (see the remark after Definition 7.7). Since $D\in[a,B]$, we have that $(r_{n}(D),B)\in\mathcal{D}$ for some $n$, as desired. ∎
@@
Recall that for an uncountable cardinal $\kappa$, the forcing poset $\operatorname{Coll}(\omega,<\!\kappa)$ is the Lévy collapse, which collapses $\kappa$ to $\omega_{1}$. We shall follow the proof layout presented in [15] (for the proof of Theorem 4.1). The following lemma is due to Solovay, and a proof may be found in [26]. For the rest of this section, we fix a ground model $\mathbf{V}$ of $\mathsf{ZFC}$, and all vector spaces are over $\mathbb{F}_{2}$.
@@
Let $\mathcal{H}\in\mathbf{V}[G]$ be a selective coideal in $E^{[\infty]}$, and let $\mathcal{X}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ be a subset of $E^{[\infty]}$. Since every set in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is ordinal-definable from some real in $\mathbf{V}[G]$, there exists a ternary formula $\phi$, some real $r\in\mathbf{V}[G]$ and a sequence of ordinals $\alpha$ such that for all $B\in(E^{[\infty]})^{\mathbf{V}[G]}$:
@@
We need to show that for all $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $a,B]\subseteq\mathcal{X}$ or $a,B]\subseteq\mathcal{X}^{c}$. We fix some $A\in\mathcal{H}$ and $a\in E^{[<\infty]}\mathord{\upharpoonright}A$.
Since $\kappa$ is Mahlo, we may fix some inaccessible $\lambda<\kappa$ such that $\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]\in\mathbf{V}[G\mathord{\upharpoonright}\lambda]$, and $\mathbf{V}[G\mathord{\upharpoonright}\lambda]\models\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$ is a selective coideal, and denote $\overline{\mathcal{H}}:=\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. By the $\kappa$-chain condition of $\operatorname{Coll}(\omega,<\!\kappa)$, we may also assume that $A\in\overline{\mathcal{H}}$ and $r\in\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. Let $\phi^{*}$ be the formula such that $\phi^{*}(x,y,w)$ holds iff $\Vdash_{\operatorname{Coll}(\omega,<\kappa)}\phi(\dot{x},\check{y},\check{w})$. Let:
$\overline{\mathcal{X}}:=\{B\in E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]:V[G\mathord{\upharpoonright}\lambda]\models\phi^{*}[B,r,\alpha]\}.$
@@
$B\in\mathcal{X}$ $\iff\mathbf{V}[G]\models\phi[B,r,\alpha]$
@@
Let $\mathcal{U}$ be a $(\overline{\mathcal{H}},\leq^{*})$-generic filter such that $A\in\mathcal{U}$. Note that by Proposition 7.6, $\mathcal{U}$ is a selective ultrafilter. Since $(\overline{\mathcal{H}},\leq^{*})$ adds no reals, we have that $E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]=E^{[\infty]}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda]$. This implies that $\overline{\mathcal{X}}=\mathcal{X}\cap\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Furthermore, we have that:
@@
Let $B\in E^{[\infty]}\cap\mathbf{V}[G][\mathcal{U}]$ be $\mathbb{M}_{\mathcal{U}}$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$. Then:
@@
We work in $\mathbf{V}[G][\mathcal{U}]$. Since $\kappa$ is inaccessible, $\operatorname{Coll}(\omega,<\!\lambda)*\dot{\mathbb{M}}_{\mathcal{U}}$ is a poset in $\mathbf{V}$ of size $\kappa$, so by Lemma 7.10 applied to the model $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}]$ there exists a filter $G^{*}$, $\operatorname{Coll}(\omega,<\!\kappa)$-generic over $\mathbf{V}[G\mathord{\upharpoonright}\lambda][\mathcal{U}][B]$ such that:
@@
We may now prove Corollary 1.8. Given an almost disjoint family $\mathcal{A}$ of subspaces, we let:
$\overline{\mathcal{A}}:=\{B\in E^{[\infty]}:\langle B\rangle\subseteq V\text{ for some }V\in\mathcal{A}\}.$
@@
Suppose for a contradiction that $\mathcal{A}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$ is a full mad family of subspaces. Note that $\overline{\mathcal{A}}\in\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$. By Proposition 4.16, $\mathcal{H}(\mathcal{A})$ is a selective coideal, so by Theorem 1.7 we have that $\mathcal{A}$ is $\mathcal{H}$-Ramsey. Let $B\in\mathcal{H}(\mathcal{A})$, and let $C\in\mathcal{H}(\mathcal{A})\mathord{\upharpoonright}A$ be such that $\emptyset,C]\subseteq\overline{\mathcal{A}}$ or $\emptyset,C]\subseteq\overline{\mathcal{A}}^{c}$. Since $C\in\mathcal{H}(\mathcal{A})$ and $\mathcal{H}(\mathcal{A})\cap\overline{\mathcal{A}}=\emptyset$, the first case is not possible. On the other hand, since $\mathcal{A}$ is mad, there exists some $V\in\mathcal{A}$ which is compatible with $C$, i.e. there exists some $D\in E^{[\infty]}$ such that $D\leq C$ and $D\leq A$. Then $D\in[\emptyset,C]\cap\overline{\mathcal{A}}$, another contradiction. ∎
We conclude this section by addressing the difficulty in removing the additional large cardinal hypothesis on $\kappa$. In [15], Neeman–Norwood successfully reduced the required large cardinal strength of $\kappa$ in the original Mathias’ theorem [13] from being Mahlo to just inaccessible, with the additional requirement that the selective coideal $\mathcal{H}$ lies in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$. Their approach is as follows:
  1. (1)
    For an ultrafilter $\mathcal{U}$ on $\omega$, they defined the diagonalisation forcing $\mathbb{P}_{\mathcal{U}}$, and show that $\mathbb{P}_{\mathcal{U}}$ satisfies the Prikry and Mathias properties (where $\mathcal{U}$ need not be selective).
  2. (2)
    Let $\operatorname{Coll}(\omega,<\!\kappa)$ be the Lévy collapse of an inaccessible cardinal, and let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. There exists some $\beta<\kappa$ such that $\overline{\mathcal{H}}=\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\beta]\in\mathbf{V}[G\mathord{\upharpoonright}\beta]$. $\overline{\mathcal{H}}$ is a coideal on $\omega$, but it may not be selective. Let $\mathcal{U}\subseteq\overline{\mathcal{H}}$ be any ultrafilter on $\omega$ in $\mathbf{V}[G\mathord{\upharpoonright}\beta]$.
  3. (3)
    Repeat the argument of Mathias with Mathias forcing $\mathbb{M}_{\mathcal{U}}$ replaced by diagonalisation forcing $\mathbb{P}_{\mathcal{U}}$.
Using the definition of ultrafilter for infinite block sequences (see Definition 7.5), it is possible to define a version of diagonalisation forcing for infinite block sequences, and it does satisfy the Prikry and Mathias properties. The issue lies in the second step — if $\mathcal{H}$ is a selective coideal in $E^{[\infty]}$ in $\mathbf{V}[G]$ (or $\mathbf{L}[G]^{\mathbf{V}[G]}$), it is unclear if $\overline{\mathcal{H}}=\mathcal{H}\cap\mathbf{V}[G\mathord{\upharpoonright}\beta]$ is a coideal in $E^{[\infty]}$. It is even less clear if we may obtain an ultrafilter $\mathcal{U}\subseteq\overline{\mathcal{H}}$ in $\mathbf{V}[G\mathord{\upharpoonright}\beta]$ — the only current known construction of ultrafilters in $E^{[\infty]}$ assumes set theoretic hypotheses such as $\mathsf{CH}$ or $\mathsf{MA}$, and define an ultrafilter as the $\leq$-upward closure of a $\leq^{*}$-decreasing chain of block sequences of size continuum (see Theorem 5.3 of [23] and [3]). This requires one to repeatedly obtain weak diagonalisations (as in Definition 3.6) at countable limit ordinal stages of the induction. However, as $\overline{\mathcal{H}}$ need not satisfy the (p)-property, this construction may not be performed inside $\overline{\mathcal{H}}$ to obtain an ultrafilter $\mathcal{U}\subseteq\overline{\mathcal{H}}$.
@@ 8. Further remarks and open problems
Most of the current approaches to the non-existence of mad families in $\omega]^{\omega}$ employ some Ramsey theoretic techniques (such as [13] [15] [20]). Consequently, it is difficult to use these approaches to prove the non-existence of mad families of block subspaces over $\mathbb{F}\neq\mathbb{F}_{2}$. Perhaps the approach highlighted in [28] may be used to answer the following question:
@@
Let $\kappa$ be an inaccessible cardinal, and let $G$ be $\operatorname{Coll}(\omega,<\!\kappa)$-generic. Are there mad families of subspaces (over any countable field) in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$? What if $\kappa$ is Mahlo?